John is 12 years older than his brother Peter. Four years ago John was three times as old as Peter was
then. Suppose Peter is now x years old. Select the statement that is not true.

Answer by macston(5194)

.
P=Peter's age; J=John's age=P+12
.
J-4=3(P-4)
(P+12)-4=3P-12
P+8=3P-12
20=2P
10=P
Peter is 10 years old.
x=10 years
.
J=P+12=10+12=22
John is 22 years old.
.
No statements were given to select.